OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#10798 — Optical loop Paris
Incident Report for Web Cloud
Resolved
There is a cut in the fibre on the optical loop ANJ75, BAS75, BDV75, BEA75, GUT75, PRO75, TUI75, VOL75.


The incident has been declared to the fibre operator.

Security plays its role and the service is not impacted.

Update(s):

Date: 2014-05-16 23:40:01 UTC
Security is operational again.

Date: 2014-05-16 15:16:32 UTC
The sites are available again.

A cut was detected 900m from PRO75 to GUT75.

A welding team has been sent out to fix the issue.

Date: 2014-05-16 14:35:39 UTC
The teams are intervening on the PRO75 and GUT75 sites.
They have been isolated.

Date: 2014-05-16 14:26:59 UTC
The ANJ75 and GUT75 DSLAMs are available again since 15:45 CET.
We're working to restore the security of the optical loop.

Date: 2014-05-16 14:25:23 UTC
The ANJ75 and GUT75 DSLAMs are no longer available.

We will investigate.
Posted May 16, 2014 - 09:07 UTC